OpenStudy (anonymous):

X+6 and x-6 is equal to 48 Solve for x

OpenStudy (tkhunny):

?? Please supply the whole problem statement. That can't be it.

OpenStudy (anonymous):

The perimeter of a Rectangle with sides Is x+6 and x-6 equal to 48 inches. What is the value of x

OpenStudy (tkhunny):

Perimeter. If the sides are 5 and 6, what is the perimeter?

OpenStudy (jkristia):

To find the perimeter, add the 4 sides, you know side A is x+6 and side B = x-6 so 2A + 2B = 48 2(x+6) + 2(x-6) = 48 and go from there

OpenStudy (anonymous):

The answer is 12...right

OpenStudy (tkhunny):

Go from where? You did the hard part - the part the student was supposed to be learning!